### v1

@@ IV A classification scheme
This is however not the general situation. A physically interesting counter-example is that of identical particles SSD ; DowkerAustin  [^7] Quantization of a system of identical particles has been considered first in JMS67 , and then in Laidlaw . Asorey and Boya AsoreyBoya studied the case $\pi_{1}\simeq{\mathds{Z}}_{p}$ for $p>1$..
@@
where $z$ is the interchange of two configurations, $(q_{1},q_{2})\to(q_{2},q_{1})$. Thus we have two prequantum lifts of $\pi_{1}$ and thus two prequantisations, the first of which is trivial, while the second is twisted. The first one corresponds to bosons, the second one to fermions $\chi_{2}$ is not of the form (IV.1)  [^8] These statements are valid in at least 3 space dimensions. In the plane, the two-particle homotopy is richer and leads to anyons Leinaas77 ; Wilczek82 ..
@@ Acknowledgements.
Note added in 2024. This note was originally published in a conference proceedings HPClausthal . During the long years after its publication I came across several important papers related to the subject but not included in the original, rather succinct note, adressed to specialists in differential geometric methods of physics. In order to put it into wider perspectives and make it more readable to physicists, I decided to make available online. While following as much as I could the original note, I added many commentaries and explanations (in colored footnotes) as well as a considerably extended reference list [12 - 29]. I have also trasferred to an Appendix certain geometric details which do not belong to the standard toolbox of physicists.

### v2

@@ IV A classification scheme
This is however not the general situation. A physically interesting counter-example is that of identical particles SSD ; DowkerAustin  [^7] Quantization of a system of identical particles has been considered in JMS67 ; Laidlaw . Asorey and Boya AsoreyBoya studied the case $\pi_{1}\simeq{\mathds{Z}}_{p}$ for $p>1$..
@@
where $z$ is the interchange of two configurations, $(q_{1},q_{2})\to(q_{2},q_{1})$. Thus we have two prequantum lifts of $\pi_{1}$ and thus two prequantisations, the first of which is trivial, while the second is twisted. The first one corresponds to bosons, the second one to fermions $\chi_{2}$ is not of the form (IV.1)  [^8] These statements are valid in at least 3 space dimensions. In the plane, the two-particle homotopy is larger and leads to anyons Leinaas77 ; Wilczek82 ..
@@ Acknowledgements.
Note added in 2024. This note was originally published in a conference proceedings HPClausthal . During the long years after its publication I came across several important papers related to the subject but not included in the original, rather succinct note, adressed to specialists in differential geometric methods of physics. In order to put it into wider perspectives and make it more readable to physicists, I decided to make available online. While following as much as I could the original note, I added many commentaries and explanations (in footnotes) as well as a considerably extended reference list [12 - 29]. I have also trasferred to an Appendix certain geometric details which do not belong to the standard toolbox of physicists.